OpenStudy (anonymous):

8sin^2 x + 6sin x = 9 Solve for 0 ≤ x ≤ 2π Restricted (0, 2π)

OpenStudy (anonymous):

\[\frac{ 9 }{ 2 } and \frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

It's supposed to be in radians, uhh... I know I need the quadratic formula for it, and I'm supposed to get two values to set equal to zero, then find those values on a unit circle. A similar problem might give me sin x - 1 = 0, so I'd solve that for sin x; sin x = 1, and then find where sin is equal to 1 on the unit circle, which would be my answer.

OpenStudy (anonymous):

sinx = 3/4

OpenStudy (anonymous):

Do you mind showing me how you got that answer?